Circular No. 7992

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 54P/DE VICO-SWIFT = P/2002 T4 (NEAT) A. Nakamura, Kuma, Ehime, Japan, reports that K. Muraoka (Kochi,

Japan) has identified comet P/2002 T4 (cf. IAUC 7991) with 54P, last seen in 1965. The indicated correction to the prediction on MPC 34423 (ephemeris on MPC 46016) is ΔT = −7.5 days. Calculations by B. G. Marsden, Smithsonian Astrophysical Observatory, indicate that the comet passed 0.16 AU from Jupiter on 1968 Oct. 18; his revised orbital elements from 34 observations, 1965–2002:

Epoch = 2002 July 25.0 TT

T = 2002 July 30.93980 TT  ω = 2.^{o}05527
e = 0.4306928  Ω = 358.96433 }2000.0
q = 2.1449799 AU  i = 6.08493

a = 3.7677022 AU n^{o} = 0.13476876 P = 7.313 years

V838 MONOCEROTIS R. M. Wagner, LBT Observatory; and S. G. Starrfield, Arizona State

University, report: “An optical spectrum (range 390–750 nm; resolution 0.35 nm) of V838 Mon was obtained on Sept. 25.498 UT with the 6.5-m MMT (+ blue channel spectrograph). Our spectrum is composite and thus confirms an earlier report of a weak blue continuum in low-S/N spectra (Desidera and Munari, IAUC 7982). Strong TiO and VO bands are seen in the red, in agreement with the recent spectra, suggesting a late-M-giant component. However, at shorter wavelengths, there are strong and narrow absorption lines of Hβ, Hγ, Hδ, and He I (447.1, 587.5 nm). The equivalent width of Hβ is 0.6 nm. Narrow emission lines arising from [O I] (630.0, 636.3 nm), Mg I (457.1, 516.7 nm), and many other weaker emission and absorption lines are also present. The spectrum exhibits strong interstellar absorption features of Na I D, 578.0 nm, and perhaps 443.0 nm. The equivalent width of 578.0 nm implies E(B − V) = 0.54 mag, corresponding to 1.7 mag of visual extinction and consistent with a previous estimate (Munari et al. 2002, A.Ap. 389, L51) based on outburst spectra. The composite spectrum suggests that this is a binary system consisting of a cool giant and a hot-star companion. Additional blue spectroscopy to ascertain the nature of the hot component is required.”

2002 October 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green